# Speech Robust Bench: A Robustness Benchmark For Speech Recognition

### v2

@@ 6 Conclusion
We propose SRB, a comprehensive benchmark designed to standardize the robustness evaluations of Automatic Speech Recognition (ASR) models. To the best of our knowledge, SRB is the first ASR robustness benchmark that contains long-form speech, and expands the set of distortions to include special effects, speaker attribute modification and adversarial attacks. We demonstrate the utility of SRB in evaluating the robustness of ASR models via several concrete examples, as well as its potential to facilitate evaluations of other aspects of trustworthy AI, like fairness. We believe that SRB will enable rigorous robustness evaluations of ASR models in a highly standardized manner, allowing easy comparisons between existing and new approaches. To further facilitate robustness evaluations for researchers and model developers, we release transformed test sets in English and Spanish. We anticipate that this will make robustness evaluations more prevalent and encourage model developers to consider robustness as a key metric for improvement.
@@ Appendix A Perturbation Generation/Application Procedure
We use the recordings of environmental noises from the test/eval subsets of ESC-50  [71], MS-SNSD [60], MUSAN [14] and WHAM [9]. We create a separate perturbed version of the clean data using each of these noise datasets. To do so, for each test utterance we sample a random environmental noise and add it to the audio signal at the specified SNR. We clip the noise if it is longer than the speech, and repeat it if it is shorter than the speech. We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
@@
The simulated and real RIRs from [16] are applied to clean recordings. As a measure of intensity, RT60 is estimated for the simulated RIRs using Sabine’s formula with the room dimensions and absorption coefficient provided in the dataset. For the real RIRs, we compute the SRMR [72] using the implementation from https://github.com/aliutkus/speechmetrics/tree/master. The severity is defined in increasing RT60s for the synthetic RIRs, and decreasing SRMR for the real RIRs. Table 1 shows the average RT60/SRMS in each severity level. During evaluation, a random RIR having the given severity level is sampled for each test recording.
@@
We use use YourTTS [59] from Coqui.ai [^2] https://github.com/coqui-ai/TTS to synthesize audio from textual transcripts in a given speaker’s style. The transcripts from the test clean subset of LibriSpeech are used. The target speakers are drawn from the VCTK corpus [73], which contains accented speech from 12 accents. For each transcript a random speaker is chosen to synthesize the audio.
Untargeted Adversarial Attack:
The adversarial perturbations are computed using an untargeted PGD adversarial attack. The attack is computed as follows. First, the maximum possible L2 norm of the noise is determined by solving the equation for SNR for the norm of the noise.
@@
where $\delta$ is the noise, $x$ is the audio signal and SNR is the upper bound on the SNR in the final signal. Then, we follow the approach of [43] and optimize $\delta$ using Projected Gradient Descent (PGD) to maximize the divergence between the true and predicted transcriptions.
A.1 Crosstalk and Music
We use crosstalk and music audios from MUSAN [14]. We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
A.2 Accents
We select a subset of audios from the test set of Common Voice 17. The selected audios satisfied the following criteria: (1) the speaker’s accent must be present in the metadata, (2) the accent must not be American, (3) should be clean. The last criterion is satisfied if the DNSMOS [10] score of the recording is at least 3.4. The resulting subset contains 640 recordings. The most popular accent in this set is South Asian (India, Pakistan, Sri Lanka) (2̃5%), followed by British English (2̃5%).
A.3 Inter-Personal Communications
We CHiME-6 [7] and AMI [8] to obtain recordings of people in social scenarios (dinner party and meetings). In both these datasets, the speakers are recorded through lapel microphone and a room microphone resulting in near and far field recordings. We use both types of recordings and show separate results for them. We remove recordings that contain less than three words since they are often fillers.
Untargeted Utterance Agnostic Adversarial Attack:
We use the method of [52] to compute utterance agnostic adversarial perturbations. Concretely, we compute a single perturbation vector that can cause the accuracy of the model on several (ideally all) inputs to degrade severely. The main difference between the universal attack and the PGD attack is that the latter computes a perturbation vector for each input, whereas the former computes a single perturbation that is expected to successfully attack any input to the model. Once we compute the perturbation we add it to the audio signal at the specified SNR.
We use the robust_speech package [74] to generate adversarial perturbations.
@@ Appendix B Additional definitions
In this work, we use SRB to conduct a fairness assessment based on robustness disparities across population subgroups (English vs. Spanish speech; male vs. female speakers). Most of the state of the art quantifies fairness in terms of predictive performance disparities. This occurs in the domain of fairness in ASR systems [63, 64, 70] and similar prevalence is found in other domains [75, 76]. However, previous work in the domain of ASR also considered robustness disparities as an alternative fairness notion [77].
@@ Appendix C Models
# Params (M) Data (hrs) WER
Language Model model (abrv)
EN canary-1b [25] cnry-1b 1,000.0 85,000 6.0
deepspeech[67] ds 86.0 960 26.5
hubert-large-ls960-ft[39] hubt-lg 317.0 60,000 8.4
hubert-xlarge-ls960-ft[39] hubt-xl 964.0 60,000 8.4
mms-1b-fl102 [78] $m$s-1b 964.6 55,000 22.8
speecht5 asr[66] spch-t5 154.6 960 22.1
wav2vec2-base-960h [38] w2v2-bs 95.0 960 11.3
wav2vec2-large-960h [38] w2v2-lg 317.0 960 9.7
wav2vec2-large-960h-lv60-self [65] w2v2-lg-slf 317.0 60,000 7.7
wav2vec2-large-robust-ft-libri-960h [58] w2v2-lg-rob 317.0 63,000 8.9
whisper-base [19] wsp-bs 74.0 680,000 9.6
whisper-base.en [19] wsp-bs.en 74.0 563,000 5.1
whisper-large-v2 [19] wsp-lg 1,550.0 680,000 8.0
whisper-medium [19] wsp-md 769.0 680,000 7.9
whisper-medium.en [19] wsp-md.en 769.0 563,000 4.1
whisper-small [19] wsp-sm 244.0 680,000 8.3
whisper-small.en [19] wsp-sm.en 244.0 563,000 4.0
whisper-tiny [19] wsp-tn 39.0 680,000 11.3
whisper-tiny.en [19] wsp-tn.en 39.0 563,000 10.1
ES canary-1b [25] cnry-1b 1,000.0 85,000 3.2
mms-1b-fl102 [78] $m$s-1b 964.6 55,000 15.7
wav2vec2-base-10k-voxpopuli-ft-es [68] w2v2-bs-es 94.4 10,116 25.7
wav2vec2-large-xlsr-53-spanish [69] w2v2-lg-es 315.4 54,350 6.8
whisper-base [19] wsp-bs 74.0 680,000 14.8
whisper-large-v2 [19] wsp-lg 1,550.0 680,000 5.8
whisper-tiny [19] wsp-tn 39.0 680,000 23.3
@@ Appendix D Fine Grained Analyses
The following figures present fine-grained analyses of robustness. These figures may be referenced by the main text but were not included in the main body in the interest of space. Figures 7, and 8 present the breakdown by perturbation of the robustness of models on English and Spanish, respectively. Figure 9 presents a breakdown of robustness by severity.

### v3

@@ 6 Conclusion
We propose SRB, a comprehensive benchmark designed to standardize the robustness evaluations of Automatic Speech Recognition (ASR) models. We demonstrate the utility of SRB in evaluating the robustness of ASR models via several concrete examples, as well as its potential to facilitate evaluations of other aspects of trustworthy AI, like fairness. We believe that SRB will enable rigorous robustness evaluations of ASR models in a highly standardized manner, allowing easy comparisons between existing and new approaches. To further facilitate robustness evaluations for researchers and model developers, we release transformed test sets in English and Spanish. We anticipate that this will make robustness evaluations more prevalent and encourage model developers to consider robustness as a key metric for improvement.
@@ Appendix A Perturbation Generation/Application Procedure
Name Subset Hours Utterances Speakers Male/Female
LibriSpeech test-clean 5.4 2620 40 20/20
TEDLIUM 3 test 3.76 1155 16 10/6
Multi-Lingual LibriSpeech (es) test 10 2385 20 10/10
CHiME-6 eval 5.25 13000 8 -
AMI test 7.35 13168 16 8/8
ESC-50 - 2.78 2000 - -
MUSAN - 108.5 2016 - -
WHAM! noise-test 9 3000 - -
MS-SNSD noise-test 0.7 51 - -
Table 4: Distributional statistics of speech (top) and noise (bottom) datasets used in SRB.
@@
We use the recordings of environmental noises from the test/eval subsets of ESC-50  (Piczak, 2015), MS-SNSD (Reddy et al., 2019), MUSAN (Snyder et al., 2015) and WHAM (Wichern et al., 2019). We create a separate perturbed version of the clean data using each of these noise datasets. To do so, for each test utterance we sample a random environmental noise and add it to the audio signal at the specified SNR. We clip the noise if it is longer than the speech, and repeat it if it is shorter than the speech. We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
@@
The simulated and real RIRs from (Ko et al., 2017) are applied to clean recordings. As a measure of intensity, RT60 is estimated for the simulated RIRs using Sabine’s formula with the room dimensions and absorption coefficient provided in the dataset. For the real RIRs, we compute the SRMR (Santos & Falk, 2014) using the implementation from https://github.com/aliutkus/speechmetrics/tree/master. The severity is defined in increasing RT60s for the synthetic RIRs, and decreasing SRMR for the real RIRs. Table 2 shows the average RT60/SRMS in each severity level. During evaluation, a random RIR having the given severity level is sampled for each test recording.
@@
We use use YourTTS (Casanova et al., 2022) from Coqui.ai [^3] https://github.com/coqui-ai/TTS to synthesize audio from textual transcripts in a given speaker’s style. The transcripts from the test clean subset of LibriSpeech are used. The target speakers are drawn from the VCTK corpus (Yamagishi et al., 2019), which contains accented speech from 12 accents. For each transcript a random speaker is chosen to synthesize the audio.
Crosstalk and Music
We use crosstalk and music audios from MUSAN (Snyder et al., 2015). We use torchaudio.function.add_noise to add the noise to the speech at a given SNR.
Accents
We select a subset of audios from the test set of Common Voice 17. The selected audios satisfied the following criteria: (1) the speaker’s accent must be present in the metadata, (2) the accent must not be American, (3) should be clean. The last criterion is satisfied if the DNSMOS (Reddy et al., 2020) score of the recording is at least 3.4. The resulting subset contains 640 recordings. The most popular accent in this set is South Asian (India, Pakistan, Sri Lanka) (2̃5%), followed by British English (2̃5%).
Inter-Personal Communications
We CHiME-6 (Barker et al., 2017) and AMI (Kraaij et al., 2005) to obtain recordings of people in social scenarios (dinner party and meetings). In both these datasets, the speakers are recorded through lapel microphone and a room microphone resulting in near and far field recordings. We use both types of recordings and show separate results for them. We remove recordings that contain less than three words since they are often fillers.
Uttarance Specific Adversarial Attack:
The utterance specific adversarial perturbations are computed using the untargeted PGD adversarial attack implemented in robust_speech package (Olivier & Raj, 2022). The attack is computed as follows. First, the maximum possible L2 norm of the noise is determined by solving the equation for SNR for the norm of the noise as follows.
@@
where $\delta$ is the noise, $x$ is the audio signal and SNR is the upper bound on the SNR in the final signal. Then, we follow the approach of (Madry et al., 2018) and optimize $\delta$ using Projected Gradient Descent (PGD) to maximize the divergence between the true and predicted transcriptions. Formally stated, the attack performs the following optimization:
$\delta=\max_{\hat{\delta}:\left\lVert\hat{\delta}\right\rVert_{2}\leq\epsilon_{\text{SNR}}}L_{M}(x+\hat{\delta},y),$ (3)
where $L_{M}$ is the loss function used to train the ASR model, $M$, such as CTCLoss or NLLLoss.
Utterance Agnostic Adversarial Attack:
We use the method of (Neekhara et al., 2019), as implemented in robust_speech package (Olivier & Raj, 2022), to compute utterance agnostic adversarial perturbations. The main difference between the universal attack and the PGD attack is that the latter computes a perturbation vector for each input, whereas the former computes a single perturbation that is expected to successfully attack any input to the model.
Formally, given a ASR model, $M$, and a development speech dataset, $\mathcal{X}^{\text{dev}}$ let $\mathcal{X}^{\text{dev}}_{\delta}=\{x+\delta|x\in\mathcal{X}^{\text{dev}}\}$ be the same dataset under additive perturbation $\delta$, and let $M(\mathcal{X}^{\text{dev}})$ and $M(\mathcal{X}^{\text{dev}}_{\delta}$ be the transcripts predicted by $M$ for $\mathcal{X}^{\text{dev}}$ and $\mathcal{X}^{\text{dev}}_{\delta}$. The utterance agnostic attack uses PGD to optimize $\delta$ such that $\left\lVert\delta\right\rVert_{\infty}\leq\epsilon$ and the Character-Error Rate (CER) (see § B) between $M(\mathcal{X}^{\text{dev}})$ and $M(\mathcal{X}^{\text{dev}}_{\delta})$ is at least $t$, i.e. $\text{CER}^{M}(\mathcal{X}^{\text{dev}}_{\delta},M(\mathcal{X}^{\text{dev}}))\geq t$ (using the notation from § B). Similar to the utterance-specific attack, the value of $\epsilon$ is determined by the maximum allowable SNR using eq.(2), except that $\ell_{\infty}$ norms are used instead of $\ell_{2}$ norms. The full algorithm is described in Algorith 1.
Algorithm 1 Utterance Agnostic Attack Algorithm
Speech data $\mathcal{X}^{\text{dev}}$ and , ASR model $M$ and its loss function, $L_{M}$ (e.g. CTCLoss, NLLLoss, etc.), allowed SNR $s$, learning rate, $\alpha$, max epochs $e_{max}$, max iterations per sample $i_{max}$, target attack success rate, $t_{sr}$, target Character-Error Rate (CER) (see § B), $t_{cer}$
procedure CER($a,b$)
  return $\text{EditDistance}(a,b)/len(b)$ $\triangleright$ $a$ and $b$ are character sequences
end procedure
procedure SuccessRate($\mathcal{X}$)
  return $\sum_{x\in\mathcal{X}}I[CER(M(x+v),M(x))>t_{cer}]$
end procedure
procedure SNRToNorm($x$, SNR)
@@
end procedure
$\epsilon\leftarrow\sum_{x\in\mathcal{X}^{\text{dev}}}SNRToNorm(x,s)/|\mathcal{X}^{\text{dev}}|$
$v\leftarrow 0$
$e\leftarrow 0$
while $SuccessRate<t_{sr}$ and $e<e_{max}$ do
  for $(x,y)\in\mathcal{X}^{\text{dev}}$ do $\triangleright$ $x$ is the audio, $y$ is the transcript
   $i\leftarrow 0$
   $r\leftarrow 0$
   while $CER(M(x+v+r),M(x))>t_{cer}$ and $i<i_{max}$ do
     $\Delta_{r}\leftarrow\alpha sign(\nabla_{r}0.5\left\lVert r\right\rVert_{2}-L_{M}(x+v+r,y))$
     $r\leftarrow clip_{\epsilon}\{r-\Delta_{r}+v\}-v$
     $i\leftarrow i+1$
   end while
   $v\leftarrow clip_{\epsilon}\{r+v\}$
  end for
  $e\leftarrow e+1$
end while
Once we compute the perturbation we add it to the test audios ($\mathcal{X}^{\text{test}}$) at the specified SNR using torchaudio.function.add_noise. For LibriSpeech, we use 500 utterances from test-dev as $\mathcal{X}^{\text{dev}}$ and test-clean as $\mathcal{X}^{\text{dev}}$. For TEDLIUM, we use the full dev and test sets as $\mathcal{X}^{\text{dev}}$ and $\mathcal{X}^{\text{test}}$. For Multi-Lingual LibriSpeech, we use 500 utterances from the dev set in the relevant language as $\mathcal{X}^{\text{dev}}$ and the full test set of the same language as $\mathcal{X}^{\text{test}}$.
@@ Appendix B Additional definitions
The Character Error Rate (CER) can also be defined similarly, by using the character-level edit distance in the above equation.
@@
In this work, we use SRB to conduct a fairness assessment based on robustness disparities across population subgroups (English vs. Spanish speech; male vs. female speakers). Most of the state of the art quantifies fairness in terms of predictive performance disparities. This occurs in the domain of fairness in ASR systems (Liu et al., 2022; Veliche & Fung, 2023; Koenecke et al., 2020) and similar prevalence is found in other domains (Julia Angwin, ; Solans Noguero et al., 2023). However, previous work in the domain of ASR also considered robustness disparities as an alternative fairness notion (Nanda et al., 2021).
@@ Appendix C Models
Lang Model Abrv. Params (M) Data (hrs) WER
EN canary-1b (NVIDIA, ) cnry-1b 1,000.0 85,000 6.0
deepspeech(Amodei et al., 2016) ds 86.0 960 26.5
hubert-large-ls960-ft(Hsu et al., 2021a) hubt-lg 317.0 60,000 8.4
hubert-xlarge-ls960-ft(Hsu et al., 2021a) hubt-xl 964.0 60,000 8.4
mms-1b-fl102 (Pratap et al., 2023) $m$s-1b 964.6 55,000 22.8
speecht5 asr(Ao et al., 2022) spch-t5 154.6 960 22.1
wav2vec2-base-960h (Baevski et al., 2020) w2v2-bs 95.0 960 11.3
wav2vec2-large-960h (Baevski et al., 2020) w2v2-lg 317.0 960 9.7
wav2vec2-large-960h-lv60-self (Xu et al., 2021) w2v2-lg-slf 317.0 60,000 7.7
wav2vec2-large-robust-ft-libri-960h (Hsu et al., 2021b) w2v2-lg-rob 317.0 63,000 8.9
whisper-base (Radford et al., 2023) wsp-bs 74.0 680,000 9.6
whisper-base.en (Radford et al., 2023) wsp-bs.en 74.0 563,000 5.1
whisper-large-v2 (Radford et al., 2023) wsp-lg 1,550.0 680,000 8.0
whisper-medium (Radford et al., 2023) wsp-md 769.0 680,000 7.9
whisper-medium.en (Radford et al., 2023) wsp-md.en 769.0 563,000 4.1
whisper-small (Radford et al., 2023) wsp-sm 244.0 680,000 8.3
whisper-small.en (Radford et al., 2023) wsp-sm.en 244.0 563,000 4.0
whisper-tiny (Radford et al., 2023) wsp-tn 39.0 680,000 11.3
whisper-tiny.en (Radford et al., 2023) wsp-tn.en 39.0 563,000 10.1
ES canary-1b (NVIDIA, ) cnry-1b 1,000.0 85,000 3.2
mms-1b-fl102 (Pratap et al., 2023) $m$s-1b 964.6 55,000 15.7
wav2vec2-base-10k-voxpopuli-ft-es (Wang et al., 2021b) w2v2-bs-es 94.4 10,116 25.7
wav2vec2-large-xlsr-53-spanish (Conneau et al., 2020) w2v2-lg-es 315.4 54,350 6.8
whisper-base (Radford et al., 2023) wsp-bs 74.0 680,000 14.8
whisper-large-v2 (Radford et al., 2023) wsp-lg 1,550.0 680,000 5.8
whisper-tiny (Radford et al., 2023) wsp-tn 39.0 680,000 23.3
@@ Appendix D Fine Grained Analyses
The following figures present fine-grained analyses of robustness. These figures may be referenced by the main text but were not included in the main body in the interest of space. Figure 8 provides an overview of the accuracy and robustness of the various models. Figures 9, and 10 present the breakdown by perturbation of the robustness of models on English and Spanish, respectively. Figure 11 presents a breakdown of robustness by severity.
@@ Appendix F Dataset Licenses
Table 7: Accuracy and robustness of ASR models on all datasets
category clean accent audio proc noise (env) noise (white) sFX social (FF) social (NF) spatial synth speech AVG Adv (UA) Adv (US) AVG
lang dataset model (abrv) WER NWERD WERD
du MLS-DU cnry-1b 4.0 - 15.5 5.3 6.0 12.4 - - 12.2 - 10.3 - - -
wsp-lg 7.3 - 14.3 2.6 6.5 4.1 - - 5.3 - 6.6 - - -
$m$s-1b 14.6 - 21.6 8.1 12.0 14.3 - - 18.9 - 15.0 - - -
wsp-bs 20.4 - 93.8 14.0 26.3 34.5 - - 24.7 - 38.6 - - -
en LibriSpeech prkt-rnnt-1.1b 1.6 - 5.1 2.6 1.8 3.7 - - 8.0 3.0 4.0 - - -
cnry-1b 1.7 - 12.3 3.0 2.7 12.7 - - 9.3 4.2 7.4 7.7 61.7 34.7
prkt-rnnt-0.6b 1.8 - 11.2 3.4 3.0 6.0 - - 9.4 3.8 6.1 - - -
w2v2-lg-slf 1.8 - 16.9 9.2 17.1 12.4 - - 14.4 5.5 12.6 12.5 58.0 35.2
prkt-ctc-0.6b 2.0 - 10.9 3.5 3.4 7.3 - - 10.4 4.7 6.7 - - -
prkt-ctc-1.1b 2.0 - 6.2 3.0 3.3 6.4 - - 9.1 4.4 5.4 - - -
hubt-xl 2.1 - 18.2 11.1 16.3 13.3 - - 14.7 5.5 13.2 26.7 54.4 40.5
hubt-lg 2.1 - 11.6 9.0 9.1 12.3 - - 15.0 6.5 10.6 34.9 50.0 42.5
sb-cnfmr 2.6 - 9.4 11.0 15.2 11.3 - - 19.0 7.4 12.2 - - -
w2v2-lg-rob 2.6 - 18.1 10.2 15.1 15.4 - - 16.9 5.9 13.6 45.8 61.0 53.4
sb-cnfmr-rnnt 2.7 - 15.7 9.8 17.2 9.6 - - 20.4 8.4 13.5 - - -
w2v2-lg 3.0 - 24.7 14.4 28.2 15.8 - - 21.8 8.6 18.9 29.0 44.3 36.6
w2v2-bs 3.7 - 30.6 18.9 36.4 20.3 - - 26.9 9.6 23.8 29.5 53.6 41.5
wsp-md 3.9 - 19.1 2.4 3.3 2.1 - - 3.8 5.3 6.0 1.7 51.9 26.8
wsp-lg 3.9 - 6.9 2.0 3.5 1.6 - - 3.0 4.0 3.5 3.4 43.6 23.5
wsp-sm.en 4.0 - 34.9 6.9 - 6.0 - - 3.9 6.2 11.6 - - -
wsp-md.en 4.1 - 16.0 4.9 - 2.8 - - 0.9 4.8 5.9 - - -
wsp-sm 4.3 - 27.9 4.0 5.4 4.5 - - 6.5 7.0 9.2 8.3 71.1 39.7
wsp-bs.en 5.1 - 48.0 12.4 - 17.9 - - 11.2 8.1 19.5 - - -
wsp-bs 5.9 - 63.6 6.6 11.0 14.8 - - 14.5 8.5 19.8 3.6 96.7 50.2
wsp-tn.en 6.4 - 57.2 8.6 14.0 25.8 - - 17.1 9.2 22.0 9.4 90.2 49.8
spch-t5 7.2 - 32.6 49.8 26.2 27.3 - - 67.9 64.3 44.7 52.4 106.4 79.4
sbcrdnn 7.2 - - 11.5 33.9 - - - - - 22.7 - - -
wsp-tn 8.2 - 68.4 13.8 17.4 26.1 - - 21.8 9.9 26.2 22.5 - 22.5
ds 15.1 - 37.9 26.2 28.9 32.0 - - 46.2 18.0 31.5 - 62.9 62.9
$m$s-1b 15.4 - 16.3 5.5 7.0 11.1 - - 11.9 8.9 10.1 6.2 44.6 25.4
TEDLIUM cnry-1b 10.2 - 15.3 3.1 2.9 10.6 - - 11.7 1.5 7.5 21.9 - 21.9
wsp-md 12.0 - 22.9 1.7 3.5 3.1 - - 4.3 0.8 6.1 5.7 61.3 33.5
wsp-lg 12.1 - 12.5 2.5 2.6 2.3 - - 4.0 0.4 4.0 9.0 63.6 36.3
wsp-sm 12.3 - 32.0 2.3 4.9 6.1 - - 5.6 1.6 8.8 2.2 69.8 36.0
wsp-bs 13.3 - 67.6 5.8 8.8 19.2 - - 8.7 4.1 19.0 1.7 80.3 41.0
w2v2-lg-slf 13.5 - 27.6 9.5 19.3 17.2 - - 18.8 9.9 17.0 1.5 24.0 12.8
wsp-tn.en 13.7 - 50.9 7.2 10.9 29.3 - - 12.9 2.8 19.0 51.7 90.7 71.2
wsp-tn 14.4 - 59.7 11.5 15.8 29.9 - - 16.2 4.3 22.9 1.3 110.9 56.1
hubt-lg 14.7 - 19.9 9.2 12.1 16.2 - - 17.5 9.1 14.0 0.7 13.9 7.3
hubt-xl 14.7 - 24.6 11.1 17.3 16.9 - - 18.2 8.2 16.0 1.2 19.1 10.2
w2v2-lg-rob 15.2 - 24.4 8.7 15.7 18.7 - - 19.4 9.2 16.0 1.0 17.9 9.4
w2v2-lg 16.5 - 31.2 12.9 24.7 18.8 - - 25.5 9.2 20.4 4.2 17.9 11.0
w2v2-bs 18.8 - 37.0 16.5 30.0 23.3 - - 32.6 9.5 24.8 7.1 19.9 13.5
$m$s-1b 30.2 - 19.0 6.2 7.3 12.1 - - 11.5 0.0 9.4 0.0 53.1 26.6
spch-t5 37.0 - 48.8 31.8 15.2 27.5 - - 48.1 21.7 32.2 31.8 90.6 61.2
ds 38.0 - 40.1 16.1 24.3 30.0 - - 48.7 7.3 27.8 15.3 51.5 33.4
ami cnry-1b - - - - - - 31.7 13.9 - - 22.8 - - -
hubt-lg - - - - - - 51.0 27.7 - - 39.4 - - -
hubt-xl - - - - - - 51.1 27.6 - - 39.4 - - -
$m$s-1b - - - - - - 71.2 55.2 - - 63.2 - - -
prkt-ctc-0.6b - - - - - - 34.5 14.5 - - 24.5 - - -
prkt-ctc-1.1b - - - - - - 31.6 13.6 - - 22.6 - - -
prkt-rnnt-0.6b - - - - - - 37.0 16.7 - - 26.9 - - -
prkt-rnnt-1.1b - - - - - - 35.9 17.0 - - 26.4 - - -
sb-cnfmr - - - - - - 63.0 35.4 - - 49.2 - - -
sb-cnfmr-rnnt - - - - - - 58.6 31.6 - - 45.1 - - -
sbcrdnn - - - - - - 91.2 - - - 91.2 - - -
spch-t5 - - - - - - 304.3 221.9 - - 263.1 - - -
w2v2-bs - - - - - - 66.0 32.9 - - 49.4 - - -
w2v2-lg - - - - - - 59.6 30.8 - - 45.2 - - -
w2v2-lg-rob - - - - - - 51.1 24.7 - - 37.9 - - -
w2v2-lg-slf - - - - - - 50.6 26.2 - - 38.4 - - -
wsp-bs - - - - - - 39.5 19.1 - - 29.3 - - -
wsp-lg - - - - - - 29.3 15.9 - - 22.6 - - -
wsp-md - - - - - - 29.2 15.8 - - 22.5 - - -
wsp-sm - - - - - - 31.4 17.0 - - 24.2 - - -
wsp-tn - - - - - - 48.9 21.2 - - 35.0 - - -
wsp-tn.en - - - - - - 41.7 19.5 - - 30.6 - - -
chime cnry-1b - - - - - - 55.6 28.8 - - 42.2 - - -
ds - - - - - - 90.6 84.0 - - 87.3 - - -
hubt-lg - - - - - - 82.1 55.2 - - 68.7 - - -
hubt-xl - - - - - - 81.1 54.6 - - 67.8 - - -
$m$s-1b - - - - - - 126.7 98.8 - - 112.7 - - -
prkt-ctc-0.6b - - - - - - 56.6 27.4 - - 42.0 - - -
prkt-ctc-1.1b - - - - - - 53.4 26.7 - - 40.0 - - -
prkt-rnnt-0.6b - - - - - - 60.5 28.3 - - 44.4 - - -
prkt-rnnt-1.1b - - - - - - 55.9 27.3 - - 41.6 - - -
sb-cnfmr - - - - - - 92.5 74.2 - - 83.3 - - -
sb-cnfmr-rnnt - - - - - - 84.3 60.5 - - 72.4 - - -
sbcrdnn - - - - - - 92.9 87.3 - - 90.1 - - -
spch-t5 - - - - - - 527.7 388.7 - - 458.2 - - -
w2v2-bs - - - - - - 88.7 64.6 - - 76.7 - - -
w2v2-lg - - - - - - 86.9 61.7 - - 74.3 - - -
w2v2-lg-rob - - - - - - 83.4 52.3 - - 67.8 - - -
w2v2-lg-slf - - - - - - 81.6 52.0 - - 66.8 - - -
wsp-bs - - - - - - 83.4 35.5 - - 59.4 - - -
wsp-lg - - - - - - 48.6 21.6 - - 35.1 - - -
wsp-md - - - - - - 50.6 22.8 - - 36.7 - - -
wsp-sm - - - - - - 59.2 24.9 - - 42.1 - - -
wsp-tn - - - - - - 98.6 46.3 - - 72.4 - - -
wsp-tn.en - - - - - - 80.4 35.2 - - 57.8 - - -
CV cnry-1b - 4.6 - - - - - - - - 4.6 - - -
ds - 93.4 - - - - - - - - 93.4 - - -
hubt-lg - 14.1 - - - - - - - - 14.1 - - -
hubt-xl - 12.9 - - - - - - - - 12.9 - - -
$m$s-1b - 18.1 - - - - - - - - 18.1 - - -
prkt-ctc-0.6b - 5.8 - - - - - - - - 5.8 - - -
prkt-ctc-1.1b - 5.6 - - - - - - - - 5.6 - - -
prkt-rnnt-0.6b - 5.1 - - - - - - - - 5.1 - - -
prkt-rnnt-1.1b - 3.8 - - - - - - - - 3.8 - - -
sb-cnfmr - 20.3 - - - - - - - - 20.3 - - -
sb-cnfmr-rnnt - 19.1 - - - - - - - - 19.1 - - -
sbcrdnn - 44.8 - - - - - - - - 44.8 - - -
spch-t5 - 89.7 - - - - - - - - 89.7 - - -
w2v2-bs - 25.4 - - - - - - - - 25.4 - - -
w2v2-lg - 20.0 - - - - - - - - 20.0 - - -
w2v2-lg-rob - 13.9 - - - - - - - - 13.9 - - -
w2v2-lg-slf - 13.5 - - - - - - - - 13.5 - - -
wsp-bs - 10.0 - - - - - - - - 10.0 - - -
wsp-lg - 3.7 - - - - - - - - 3.7 - - -
wsp-md - 4.1 - - - - - - - - 4.1 - - -
wsp-sm - 6.5 - - - - - - - - 6.5 - - -
wsp-tn - 13.9 - - - - - - - - 13.9 - - -
wsp-tn.en - 12.7 - - - - - - - - 12.7 - - -
es MLS-ES cnry-1b 3.2 - 16.4 9.6 8.2 16.0 - - 10.9 - 12.2 26.1 84.3 55.2
wsp-lg 5.8 - 14.1 7.4 5.9 4.0 - - 2.0 - 6.7 13.7 65.0 39.4
w2v2-lg-es 6.8 - 21.2 14.7 22.1 19.0 - - 26.3 - 20.6 33.9 71.0 52.4
wsp-bs 14.8 - 90.9 22.7 27.3 32.5 - - 18.0 - 38.3 19.5 159.5 89.5
$m$s-1b 15.7 - 18.5 19.7 37.1 20.0 - - 14.9 - 22.0 7.4 53.8 30.6
wsp-tn 23.3 - 124.3 45.0 52.3 57.9 - - 41.8 - 64.3 43.1 269.9 156.5
w2v2-bs-es 25.7 - 32.4 20.4 24.2 24.1 - - 32.4 - 26.7 10.0 33.8 21.9
CV:es cnry-1b - 205.0 - - - - - - - - 205.0 - - -
$m$s-1b - 7.7 - - - - - - - - 7.7 - - -
w2v2-bs-es - 13.3 - - - - - - - - 13.3 - - -
w2v2-lg-es - 9.3 - - - - - - - - 9.3 - - -
wsp-bs - 18.2 - - - - - - - - 18.2 - - -
wsp-lg - 2.0 - - - - - - - - 2.0 - - -
wsp-tn - 30.6 - - - - - - - - 30.6 - - -
fr MLS-FR cnry-1b 6.1 - 15.2 5.2 7.3 13.0 - - 10.0 - 10.1 - - -
wsp-lg 7.7 - 15.5 3.5 8.3 5.6 - - 5.7 - 7.7 - - -
$m$s-1b 23.6 - 21.0 6.9 12.5 12.9 - - 15.6 - 13.8 - - -
wsp-bs 26.0 - 124.8 16.7 44.4 47.9 - - 25.9 - 51.9 - - -